OpenStudy (anonymous):

how can we solve log base 64 of x+log base8 of x+log base 2 of x=18?

OpenStudy (anonymous):

\[\log_{64} X+\log_{8}X +\log_{2}X =18\] is that right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok ,so do u know from where do we start ?

OpenStudy (anonymous):

no idea

OpenStudy (anonymous):

ok first lets start with the rule with is : \[\log_{a}X = \log_{\sqrt{a}} \sqrt{X} = \log_{a^2} X^2\]

OpenStudy (anonymous):

do u know this rule ?

OpenStudy (anonymous):

yeah saw it

OpenStudy (anonymous):

Do u know now from where we can start ?

OpenStudy (anonymous):

change the base or something?

OpenStudy (anonymous):

thats right ,In order to make all the bases the same so we can collect them .. lets start with \[\log_{64}X \] change it to the base 2 ,can u do it ?

OpenStudy (anonymous):

2^6?

OpenStudy (anonymous):

thats awesome right but we will use the root in order to reduce it so it will be\[\log_{64}X = \log_{\sqrt[6]{64}} \sqrt[6]{X}=\log_{2} \sqrt[6]{X}\]

OpenStudy (anonymous):

you got it ?

OpenStudy (anonymous):

8?

OpenStudy (anonymous):

lets move to \[\log_{8}X \]---> what can u do here ?

OpenStudy (anonymous):

3/2 0r 2^3?

OpenStudy (anonymous):

So we will use the cubic root in order to reduce it to two so it will be \[\log_{8}X =\log_{\sqrt[3]{8}} \sqrt[3]{X} = \log_{2} \sqrt[3]{?}\]

OpenStudy (anonymous):

sry instead of ? put X

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

no we got all the bases equal ,so what rule we gonna use now ?

OpenStudy (anonymous):

can we add them?

OpenStudy (anonymous):

now *

OpenStudy (anonymous):

we will use the rule\[\log_{a} M + \log_{a} N = \log_{a} M \times N\] ,Do u know this rule >?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

good so lets now collect them ,but first do u know how to convert a root into a power ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

are you there?

OpenStudy (anonymous):

still need help?

OpenStudy (anonymous):

yes